Unknown acid concentration in 50.0 mL H2SO4 takes 14.2 mL of 0.103 M NaOH to reach eq. point?
Answer #1

At neutralization point

mmol of H+ = mmol of OH-

M1×n×V1 = M2V2

M1 ×2×50 =0.103×14.2

M1 = 0.0146 M